2 peicent for one yes money can Dr. Ruth is gcing to borrow 37,200 bo help write s book The loan is for one year and the money can ether be borrowed st the prime rate or the LIBOR rate. Assume the prime rete is 12 percent and LIBOR 2.5 percent less. Also ssume there wil be a 345 transacion fee with LIBOR(h amount must be edded to to 2 decimal places) the imerest cost with LIBOR) a. What is the effectve interest rate on the LIBOR loanause a 360-day year, Do not round intermediate calculations, Input your answer as a percent rounded t Efective inberest rate nput your oannUse a recsve b. Which loan has the lower efecove imerest cost? o LIBOR Prime17
Answer #1

Solution :

Prime rate loan - 12%

LIBOR rate loan

Effective rate of interest = ( 9.5 % * $7,200 ) + $45 = $684 + $45 = $729

= $729 / $7200 = 10.125% ~ 10.13%

b) LIBOR

LIBOR rate is cheaper than prime rate . ( 10.13% < 12%)
Answer #2

SOLUTION :


a.


LIBOR LOAN :


Interest rate = 12 - 2.5 = 9.5 % = 0.095

Loan is for one year.

=> Loan interest cost = 7200 * 0.095 + 45 = 729 ($)



Effective rate


= 729/7200

= 0.1013

= 10.13 % (ANSWER).



 


b.


Effective interest cost on prime rate = 7200*0.12 = 864 ($)


Effective interest cost on LIBOR loan = 729 ($) 


LIBOR loan has the lower effective interest cost. (ANSWER)



answered by: Tulsiram Garg